OpenStudy (anonymous):

Sam decides to build a square garden. If the area of the garden is 4x2 + 28x + 49 square feet, what is the length of one side of the garden? (5 points) (2x + 7) feet (7x + 2) feet (2x − 7) feet (7x − 2) feet

OpenStudy (anonymous):

\[\sqrt{4 x^2+28 x+49}=\sqrt{(2 x+7)^2}=2 x+7 \]

OpenStudy (anonymous):

So does that mean that it is a?

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

Yes, "A", (2x + 7) feet